OpenStudy (anonymous):

the equation tan^x+1=sec^x

OpenStudy (anonymous):

true or false?

Nnesha (nnesha):

do you mean \[\tan^2(x)+1=\sec^{2}(x)\] ?

Nnesha (nnesha):

tan squared right ?? not tan^x

Nnesha (nnesha):

if yes then rewrite tan in terms of cos sin reciprocals : \[\tan \theta =\frac{\sin \theta}{\cos \theta}~~~~~\sin \theta=\frac{1}{\csc \theta}~~~~~~\cos \theta=\frac{1}{\sec \theta}\] the reciprocal of tan theta =1/cot theta in terms of sin cos you can write it as sin/cos which is same as 1/cot

OpenStudy (anonymous):

yes it is squared. I got false am icorrect?

OpenStudy (anonymous):

@nnesha????

OpenStudy (triciaal):

as given above by @Nnesha tan^2(x) + 1 = sec^2(x)

OpenStudy (triciaal):

it is true use that sin^2 + cos^2 = 1 and tan = sin/cos

Nnesha (nnesha):

copy these down n ur notebook http://www.purplemath.com/modules/idents.htm identities

Nnesha (nnesha):

tan theta =sin theta/ cos \[\rm \frac{ \sin^2 \theta}{ \cos^2 \theta}+1 =\sec^2 theta \] solve left side find common denominator \[\frac{ \sin^2 \theta +\cos^2 \theta }{ \cos^2 \theta } =\sec^2 \theta \] now use the special identity and the fact `1/cosx=secx`